OpenStudy (anonymous):

if tan@=a/b,find the value of sin2@,cos2@and tan2@.(@=theta.)

OpenStudy (anonymous):

what

OpenStudy (anonymous):

since tan@ = sin@/cos@, sin@ = a and cos@ = b. sin2@ = 2sin@cos@ (double angle identity), so sin2@ = 2ab. cos2@ = (cos@)^2 - (sin@)^2 = b - a

OpenStudy (anonymous):

sin2@=2ab(a^2+b^2)^-1 cos2@=(b-a)(a^2+b^2)^-1 tan2@=2ab(b-a)^-1

myininaya (myininaya):

tan@=opposite/adjacent=a/b so the hypothenues is sqrt{a^2+b^2} sin@=opp/hyp=a/(sqrt{a^2+b^2}) cos@=adj/hyp=b/(sqrt{a^2+b^2}) so sin2@=2cos@sin@=2ab/(a^2+b^2) cos2@=cos^2@-sin^2@=b^2/(a^2+b^2)-a^2/(a^2+b^2)=(b^2-a^2)/(a^2+b^2) tan2@=sin2@/cos2@=2ab/(b^2-a^2)

myininaya (myininaya):

bottom problem this should look better

myininaya (myininaya):

let me know if you have any questions shruti

OpenStudy (anonymous):

\[\tan2@ = (2\tan@)/ (1-\tan ^{2}@)\] \[= (2\sin@/\cos@) * 1/( 1- (\sin ^{2}/\cos ^{2})\] \[=2sinx/cosx * (\cos ^{2}/(\cos ^{2}/\sin ^{2}))\] \[=2sinxcosx/ (\cos ^{2} - \sin ^{2})\] \[= 2ab / (b ^{2}-a ^{2})\]

myininaya (myininaya):

remember tan means oppsite/adjacent cos means adjacent/hyp. sin means opposite/hyp.

OpenStudy (anonymous):

thx

myininaya (myininaya):

there is 2 problems on that attachment to save paper i just wrote your problem on the bottom k?

myininaya (myininaya):

if you like my answer click good answer :)